# The Symmetry Taco: Equivalences between Gapped, Gapless, and Mixed-State SPTs

### v1

@@ II Review: Anyon Condensation and the SymTFT
We begin by reviewing the basic features and some elementary examples of the SymTFT construction. In this paper, we will primarily focus on $1+1$d systems with global symmetries given by a finite group $G$ and will hence limit our discussion to this setting. The central conceit of SymTFT is that, for a $1+1$d theory with $G$ symmetry, the kinematic data for the symmetry can be encoded in a $2+1$d TO which is specified by the quantum double $\mathcal{D}(G)$ of $G$ (i.e., a discrete $G$ gauge theory) and is defined on a strip geometry. It is in this context that the TO is referred to as the SymTFT. The symmetry data of the $1+1$d theory is encoded in the SymTFT as follows: the symmetry operators, local charged operators, and their commutation relations, are related to the “magnetic” anyons, “electric” anyons, and braiding statistics of the $G$ gauge theory, respectively. Since the $1+1$d system can be thought of as living on the boundary of the $2+1$d gauge theory (in a manner that we make more precise shortly), we will often refer to the gauge theory as the “bulk” and the one-dimensional system as the “boundary.” In the following, we elaborate on the details of the SymTFT construction by focusing on the classification of gapped phases, providing explicit examples for Abelian symmetries, and reviewing how the gauging of symmetries is incorporated into this framework.
@@ II.1 The SymTFT Construction
As outlined above, the central object in the SymTFT for $1+1$d, $G$-symmetric systems is the quantum double $\mathcal{D}(G)$. The topologically distinct superselection sectors (anyons) of this $2+1$d TO (or $G$ gauge theory) are classified by the irreducible representations of $\mathcal{D}(G)$, which is the quantum double of the group algebra $\mathbb{C}[G]$ [82]. An anyon $a=([g],\pi)$ is determined by a conjugacy class $[g]$ and an irreducible representation (irrep) of the centralizer of $g$, which can be chosen to be any representative of the conjugacy class. Anyons of the form $a=([e],\pi)$ are called (electric) charges, while anyons of the form $a=([g],1)$ are called (magnetic) fluxes. All other composite anyons are referred to as “dyons”.
@@
The mathematical correspondence between gapped boundaries of $2+1$d TO and gapped $1+1$d phases can be pictorially summarized by the “sandwich” picture, where we place the bulk $\mathcal{D}(G)$ TO on a strip geometry with finite width (see Fig. 2(a)). Adopting the lexicon of Ref. [36], we denote the top boundary as the “symmetry boundary” and the bottom boundary as the “dynamical boundary”, which we also refer to as the reference and physical boundaries, respectively. Gapped boundary conditions for $2+1$d TOs are generally described via Lagrangian subalgebras, which physically correspond to maximal collections of anyons that can simultaneously condense at a boundary [^4] There are two types of anyon condensations: if the subalgebra is Lagrangian, then all anyons (except the vacuum) are confined in the resulting theory. In contrast, if the subalgebra is non-Lagrangian, then the resulting theory retains nontrivial deconfined anyons and is only partially confining.. In the SymTFT perspective, as applicable to $1+1$d $G$-symmetric phases, the reference boundary is canonically taken as the charge condensed gapped boundary, on which all anyons of the form $([e],\pi)$ are condensed. This corresponds to the Lagrangian algebra $\mathcal{L}=\text{Rep}(G)$ for a $\mathcal{D}(G)$ TO.
@@
Figure 2: (a) Thin-slab construction of SymTFT for general $1+1$d phases. (b) Illustration of bulk-boundary correspondence for symmetry operator $\prod\mathcal{X}_{i}$ and local order parameter $\mathcal{Z}_{i}^{k}$ for $\mathbb{Z}_{n}$ symmetry (shown here for $k=1$).
@@
This reference boundary encodes the kinematic data of the symmetry–the symmetry operators and the local charged operators. The symmetry operators correspond to the line operators for the confined anyons, which form the fusion category $\text{Vec}(G)$. Acting with $g\in G$ on the $1+1$d system amounts to inserting a $g$ line operator oriented parallel to the reference boundary. Since the corresponding anyons are not condensed on the reference boundary (and hence cannot be absorbed by it), such line operators act nontrivially on the dynamical boundary. In contrast, the local charged operators correspond to the anyons condensed on the reference boundary and are given by line operators which begin at the reference boundary and traverse the finite width of the slab. Such an operator can be thought of as one that tunnels an anyon $([e],\pi)$ from the reference boundary to the dynamical boundary. Since the collection of these line operators is given by the condensed anyons at the reference boundary, the local charged operators are specified by the category $\text{Rep}(G)$, where the action of the symmetry on a local charged operator is given by the action of an element $g\in G$ on the irrep $\pi$.
@@
As suggested by the name, the dynamical boundary encodes the manner in which the symmetry is dynamically realized on the $1+1$d theory–namely, whether it is preserved or is spontaneously broken. In the remainder of this Section, we will focus on the case where the data of the dynamical boundary corresponds to gapped $G$-symmetric $1+1$d phases. In this case, the dynamical boundary is fully characterized by the Lagrangian subalgebra of condensed anyons: for $\mathcal{D}(G)$ TO, these are classified by the pair $(H,[\omega])$, where $H$ is some (not necessarily normal) subgroup $H\in G$ and $[\omega]\in H^{2}(H,U(1))$ is a 2-cocycle class which describes the SPT class of the residual unbroken subgroup $H$ [77]. This data precisely coincides with the classification of $1+1$d $G$-symmetric gapped phases. Thus, for each choice of gapped dynamical boundary, the sandwich/slab construction provides a map for identifying the universal topological data characterizing the phase–the ground state degeneracy and the (string) order parameters. We now illustrate how this works in practice via simple, yet instructive, examples where $G$ is taken to be Abelian.
@@ II.2 Explicit Construction for Abelian $G$ Symmetries
For $1+1$d gapped phases with Abelian $G$ symmetry, the anyons in the corresponding bulk TO form an Abelian group under fusion, significantly simplifying the analysis (see Appendix A for details on the non-Abelian case). In particular, all anyons $a$ in this case have quantum dimension $d_{a}=1$ and can be obtained as fusion products of pure charges and pure fluxes: specifying an anyon $a=(g,\pi)$ simply amounts to specifying a charge $\pi$ and a flux $g$. Since both the group of fluxes and the group of charges are isomorphic to $G$, the group formed by the anyons under fusion is $G\times G$. The topological spin $\theta(a)$ of an anyon $a=(g,\pi)$ is given by
@@
  • All anyons in $\mathcal{L}$ have trivial mutual braiding, i.e., ,~$\forall\,\ell,\tilde{\ell}\in\mathcal{L}\,,\mathcal{B}_{\theta}(\ell,\tilde{\ell})=1$.
@@
For Abelian anyon theories, the Lagrangian subgroup $\mathcal{L}$ of condensed anyons completely determines the data of the gapped boundary [^5] Note that when $G$ is non-Abelian, the full data of a gapped boundary $(H,[\omega])$ cannot be completely recovered by the set of condensed anyons [83, 84], but this set nevertheless provides valuable intuition regarding the gapped boundary. In Appendix A, we compute which anyons from a bulk non-Abelian TO $\mathcal{D}(G)$ are condensed at the boundary labeled by $(H,[\omega])$.. In particular, the data $H\subset G$ and the $2$-cocycle class $[\omega]$ can be obtained from the Lagrangian subgroup by viewing $\mathcal{L}\subset G\times\hat{G}$ as a subgroup of anyons, where the first and second factors correspond to the groups of fluxes and charges, respectively. Then, the subgroup $H\subset G$ is the image of the projection of $\mathcal{L}$ onto the first factor. Physically, $H$ corresponds to the subgroup of fluxes that appear in the Lagrangian subgroup of condensed dyons.
@@
The first condition (antisymmetry) follows directly from the definition of $\Omega$, while the second and third properties (bilinearity) follow from $H$ being Abelian and $\gamma$ being a $2$-cocycle. The second and third conditions imply that, for every $h\in H$, the function $\Omega_{h}:h^{\prime}\mapsto\Omega(h,h^{\prime})$ is an irreducible representation of $H$, and the map $h\mapsto\Omega_{h}$ is a group homomorphism. For Abelian groups, the function $\Omega(g,h)$ satisfying the above properties completely determines the class $[\gamma]\in H^{2}(H,U(1))$.
@@
We are now in a position to illustrate the SymTFT correspondence for Abelian $G$ symmetries through examples. For each example, we will describe the data of the gapped dynamical boundary of the SymTFT in terms of both the subgroup $H$ and the $2$-cocycle class $[\omega]$, as well as via the corresponding Lagrangian subgroup of condensed anyons. From the data of the condensed anyons, we will describe how to construct a stabilizer Hamiltonian which realizes the corresponding $G$-symmetric $1+1$d gapped phase (in the non-Abelian case, these will be commuting projector Hamiltonians, which is guaranteed by the condensable algebra being Lagrangian [85]). Since these examples and the construction of the stabilizer Hamiltonians will inform our subsequent analysis of mixed-state SPTs, we would encourage even SymTFT experts to peruse the remainder of this Section.
@@
Given the bulk SymTFT, one can, in principle, construct an effective lattice Hamiltonian for the boundary which is in correspondence with the set of condensed anyons on the dynamical boundary (see e.g., Refs. [86, 24, 28, 87, 88, 41].) We will not reconstruct these arguments here but, rather, will demonstrate how the anyon data of the SymTFT can be used to construct such a Hamiltonian. In the present case, since the bulk TO is given by $\mathcal{D}(\mathbb{Z}_{n})$, we choose the boundary degrees of freedom to be $n$-qudits, with generalized Pauli operators $\mathcal{X}_{j},\mathcal{Z}_{j}$ acting on each site. These operators satisfy the $\mathbb{Z}_{n}$ Pauli algebra, $\mathcal{X}_{j}^{n}=\mathcal{Z}_{j}^{n}=1$ and $\mathcal{Z}_{j}\mathcal{X}_{j}=e^{2\pi i/n}\mathcal{X}_{j}\mathcal{Z}_{j}$. As noted in Section II.1 and depicted in Fig. 2(b), the $\mathbb{Z}_{n}$ symmetry is generated by the magnetic (or $m$) Wilson line–on the edge, we can choose it to be generated by $\mathcal{X}\equiv\prod_{i}\mathcal{X}_{i}$. A boundary on which $\langle\mathcal{X}_{i}\rangle=+1$ would then be understood as one on which $m$ is condensed. Similarly, the $\mathbb{Z}_{n}$ order parameter is given by the $e$ line stretching from the reference boundary, where it is condensed, to the dynamical boundary. Since $e$ braids with a phase of $e^{2\pi i/n}$ with $m$, it is natural to associate the action of the $e$ line terminating on the dynamical boundary with the operator $\mathcal{Z}_{i}$, since the commutation relation $\mathcal{Z}_{i}\mathcal{X}=e^{2\pi i/n}\mathcal{X}\mathcal{Z}_{i}$ matches the bulk anyon braiding. A boundary corresponding to condensation of $e$ will thus exhibit long-range order in $\mathcal{Z}_{i}$; for instance, a state with $\langle\mathcal{Z}_{i}^{\dagger}\mathcal{Z}_{j}\rangle=1$ would correspond to such an $e$-condensed boundary. This latter operator may be understood as an $e$ line emanating from a point on the dynamical boundary and terminating at an adjacent point on the same boundary, indicating that an $e$-$\overline{e}$ anyon pair can emerge from the bulk and condense on the boundary. This motivates the correspondence
@@
Exploiting this correspondence, we can write down a stabilizer Hamiltonian whose ground state subspace realizes a $1+1$d gapped phase with the symmetry data encoded in the bulk SymTFT:
[…]
@@
Next, let us consider $G=\mathbb{Z}_{2}^{A}\times\mathbb{Z}_{2}^{B}$. The bulk $\mathcal{D}(G)$ TO admits a total of six gapped boundaries to vacuum, of which five correspond to picking subgroups of $G$ but only admit a trivial $2$-cocyle (these gapped boundaries thus all correspond to $1+1$d phases where the $G$ symmetry is spontaneously broken). The remaining gapped boundary, which is obtained by picking the subgroup $H=\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ (thus there is no spontaneous symmetry breaking) along with a nontrivial cocycle, realizes the nontrivial $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SPT phase. The corresponding Lagrangian subgroup in this case is generated by
@@
as realizing the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ $1+1$d SPT. This Hamiltonian is indeed the familiar cluster-state Hamiltonian [89, 90]. From Eq. (9), we see that the two terms in this Hamiltonian can be interpreted as corresponding to the condensation of $e_{A}m_{B}$ and $e_{B}m_{A}$, respectively, on the dynamical boundary of the bulk SymTFT. Note that, in the second term we chose $X_{i+1}^{A}$ rather than $X_{i}^{A}$ in order to ensure that the stabilizers commute; this choice is not dictated by the heuristic correspondence of Eq. (9). In the ground state, the stabilizers all have eigenvalue $+1$. The string order parameters characterizing this SPT phase are
@@
where the $\ldots$ denote products of $X_{A}$ ($X_{B}$) from sites $i$ to $j-1$ (sites $i+1$ to $j$) in the first (second) line.
@@
Finally, in anticipation of our analysis for mixed-state SPTs, we consider $G=\mathbb{Z}_{4}^{A}\times\mathbb{Z}_{4}^{B}$ and focus on a particular $1+1$d phase exhibiting both symmetry breaking and SPT order under the residual symmetry. We choose $H=\mathbb{Z}_{4}^{AB}\times\mathbb{Z}_{2}^{A}$ to be the subgroup generated by the diagonal $\mathbb{Z}_{4}$ action and $\mathbb{Z}_{2}^{A}\subset\mathbb{Z}_{4}^{A}$ acting on the first factor [^7] This symmetry structure will reappear in our analysis of mixed-state examples, with the “$AB$” and “$A$” symmetries corresponding, respectively, to weak and strong symmetries. See Sec. V.. We choose the Lagrangian subgroup
@@
$H_{1d}=-\sum_{i}\Big{(}&\mathcal{Z}_{A,i-1}^{2}\mathcal{X}_{A,i-1}^{2}\mathcal{Z}_{A,i-1}^{2}+\mathcal{Z}_{B,i-1}^{2}\mathcal{X}_{B,i-1}^{2}\mathcal{Z}_{B,i}^{2}\\ +&\mathcal{Z}^{2}_{A,i-1}\mathcal{Z}^{2}_{A,i}\mathcal{Z}^{2}_{B,i-1}\mathcal{Z}^{2}_{B,i}+\mathcal{X}^{2}_{A,i}\mathcal{X}^{2}_{B,i}\\ +&\mathcal{Z}^{\dagger}_{A,i-1}\mathcal{Z}_{A,i}\mathcal{X}_{A,i}\mathcal{Z}_{B,i-1}\mathcal{Z}^{\dagger}_{B,i}\mathcal{X}_{B,i}+\text{ h.c.}\Big{)}\end{aligned}.$ (16)
@@
[…]

### v2

@@ II Review: Anyon Condensation and the SymTFT
We begin by reviewing the basic features and some elementary examples of the SymTFT construction. In this paper, we will primarily focus on 1+11+1d systems with global symmetries given by a finite group $G$ and will hence limit our discussion to this setting. The central conceit of SymTFT is that, for a 1+11+1d theory with $G$ symmetry, the kinematic data for the symmetry can be encoded in a 2+12+1d TO which is specified by the quantum double $\mathcal{D}(G)$ of $G$ (i.e., a discrete $G$ gauge theory) and is defined on a strip geometry. It is in this context that the TO is referred to as the SymTFT. The symmetry data of the 1+11+1d theory is encoded in the SymTFT as follows: the symmetry operators, local charged operators, and their commutation relations, are related to the “magnetic” anyons, “electric” anyons, and braiding statistics of the $G$ gauge theory, respectively. Since the 1+11+1d system can be thought of as living on the boundary of the 2+12+1d gauge theory (in a manner that we make more precise shortly), we will often refer to the gauge theory as the “bulk” and the one-dimensional system as the “boundary.” In the following, we elaborate on the details of the SymTFT construction by focusing on the classification of gapped phases, providing explicit examples for Abelian symmetries, and reviewing how the gauging of symmetries is incorporated into this framework.
@@ II.1 The SymTFT Construction
As outlined above, the central object in the SymTFT for 1+11+1d, $G$-symmetric systems is the quantum double $\mathcal{D}(G)$. The topologically distinct superselection sectors (anyons) of this 2+12+1d TO (or $G$ gauge theory) are classified by the irreducible representations of $\mathcal{D}(G)$, which is the quantum double of the group algebra $\mathbb{C}[G]$ [87]. An anyon $a=([g],\pi)$ is determined by a conjugacy class $[g]$ and an irreducible representation (irrep) of the centralizer of $g$, which can be chosen to be any representative of the conjugacy class. Anyons of the form $a=([e],\pi)$ are called (electric) charges, while anyons of the form $a=([g],1)$ are called (magnetic) fluxes. All other composite anyons are referred to as “dyons”.
@@
The mathematical correspondence between gapped boundaries of 2+12+1d TO and gapped 1+11+1d phases can be pictorially summarized by the “sandwich” picture, where we place the bulk $\mathcal{D}(G)$ TO on a strip geometry with finite width (see Fig. 2(a)). Adopting the lexicon of Ref. [37], we denote the top boundary as the “symmetry boundary” and the bottom boundary as the “dynamical boundary”, which we also refer to as the reference and physical boundaries, respectively. Gapped boundary conditions for 2+12+1d TOs are generally described via Lagrangian subalgebras, which physically correspond to maximal collections of anyons that can simultaneously condense at a boundary [^5] There are two types of anyon condensations: if the subalgebra is Lagrangian, then all anyons (except the vacuum) are confined in the resulting theory. In contrast, if the subalgebra is non-Lagrangian, then the resulting theory retains nontrivial deconfined anyons and is only partially confining.. In the SymTFT perspective, as applicable to 1+11+1d $G$-symmetric phases, the reference boundary is canonically taken as the charge condensed gapped boundary, on which all anyons of the form $([e],\pi)$ are condensed. This corresponds to the Lagrangian algebra $\mathcal{L}=\text{Rep}(G)$ for a $\mathcal{D}(G)$ TO.
@@
Figure 2: (a) Thin-slab construction of SymTFT for general 1+11+1d phases. (b) Illustration of bulk-boundary correspondence for symmetry operator $\prod\mathcal{X}_{i}$ and local order parameter $\mathcal{Z}_{i}^{k}$ for $\mathbb{Z}_{n}$ symmetry (shown here for $k=1$).
@@
This reference boundary encodes the kinematic data of the symmetry–the symmetry operators and the local charged operators. The symmetry operators correspond to the line operators for the confined anyons, which form the fusion category $\text{Vec}(G)$. Acting with $g\in G$ on the 1+11+1d system amounts to inserting a $g$ line operator oriented parallel to the reference boundary. Since the corresponding anyons are not condensed on the reference boundary (and hence cannot be absorbed by it), such line operators act nontrivially on the dynamical boundary. In contrast, the local charged operators correspond to the anyons condensed on the reference boundary and are given by line operators which begin at the reference boundary and traverse the finite width of the slab. Such an operator can be thought of as one that tunnels an anyon $([e],\pi)$ from the reference boundary to the dynamical boundary. Since the collection of these line operators is given by the condensed anyons at the reference boundary, the local charged operators are specified by the category $\text{Rep}(G)$, where the action of the symmetry on a local charged operator is given by the action of an element $g\in G$ on the irrep $\pi$.
@@
As suggested by the name, the dynamical boundary encodes the manner in which the symmetry is dynamically realized on the 1+11+1d theory–namely, whether it is preserved or is spontaneously broken. In the remainder of this Section, we will focus on the case where the data of the dynamical boundary corresponds to gapped $G$-symmetric 1+11+1d phases. In this case, the dynamical boundary is fully characterized by the Lagrangian subalgebra of condensed anyons: for $\mathcal{D}(G)$ TO, these are classified by the pair $(H,[\omega])$, where $H$ is some (not necessarily normal) subgroup $H\in G$ and $\omega]\in H^{2}(H,U(1))$ is a 2-cocycle class which describes the SPT class of the residual unbroken subgroup $H$ [82]. This data precisely coincides with the classification of 1+11+1d $G$-symmetric gapped phases. Thus, for each choice of gapped dynamical boundary, the sandwich/slab construction provides a map for identifying the universal topological data characterizing the phase–the ground state degeneracy and the (string) order parameters. We now illustrate how this works in practice via simple, yet instructive, examples where $G$ is taken to be Abelian.
@@ II.2 Explicit Construction for Abelian $G$ Symmetries
For 1+11+1d gapped phases with Abelian $G$ symmetry, the anyons in the corresponding bulk TO form an Abelian group under fusion, significantly simplifying the analysis (see Appendix B for details on the non-Abelian case). In particular, all anyons $a$ in this case have quantum dimension $d_{a}=1$ and can be obtained as fusion products of pure charges and pure fluxes: specifying an anyon $a=(g,\pi)$ simply amounts to specifying a charge $\pi$ and a flux $g$. Since both the group of fluxes and the group of charges are isomorphic to $G$, the group formed by the anyons under fusion is $G\times G$. The topological spin $\theta(a)$ of an anyon $a=(g,\pi)$ is given by
@@
  • All anyons in $\mathcal{L}$ have trivial mutual braiding, i.e., $\forall\,\ell,\tilde{\ell}\in\mathcal{L}\,,\mathcal{B}_{\theta}(\ell,\tilde{\ell})=1$.
@@
For Abelian anyon theories, the Lagrangian subgroup $\mathcal{L}$ of condensed anyons completely determines the data of the gapped boundary [^6] Note that when $G$ is non-Abelian, the full data of a gapped boundary $(H,[\omega])$ cannot be completely recovered by the set of condensed anyons [88, 89], but this set nevertheless provides valuable intuition regarding the gapped boundary. In Appendix B, we compute which anyons from a bulk non-Abelian TO $\mathcal{D}(G)$ are condensed at the boundary labeled by $(H,[\omega])$.. In particular, the data $H\subset G$ and the $2$-cocycle class $\omega]$ can be obtained from the Lagrangian subgroup by viewing $\mathcal{L}\subset G\times\hat{G}$ as a subgroup of anyons, where the first and second factors correspond to the groups of fluxes and charges, respectively. Then, the subgroup $H\subset G$ is the image of the projection of $\mathcal{L}$ onto the first factor. Physically, $H$ corresponds to the subgroup of fluxes that appear in the Lagrangian subgroup of condensed dyons.
@@
The first condition (antisymmetry) follows directly from the definition of $\Omega$, while the second and third properties (bilinearity) follow from $H$ being Abelian and $\gamma$ being a $2$-cocycle. The second and third conditions imply that, for every $h\in H$, the function $\Omega_{h}:h^{\prime}\mapsto\Omega(h,h^{\prime})$ is an irreducible representation of $H$, and the map $h\mapsto\Omega_{h}$ is a group homomorphism. For Abelian groups, the function $\Omega(g,h)$ satisfying the above properties completely determines the class $\gamma]\in H^{2}(H,U(1))$.
@@
We are now in a position to illustrate the SymTFT correspondence for Abelian $G$ symmetries through examples. For each example, we will describe the data of the gapped dynamical boundary of the SymTFT in terms of both the subgroup $H$ and the $2$-cocycle class $\omega]$, as well as via the corresponding Lagrangian subgroup of condensed anyons. From the data of the condensed anyons, we will describe how to construct a stabilizer Hamiltonian which realizes the corresponding $G$-symmetric 1+11+1d gapped phase (in the non-Abelian case, these will be commuting projector Hamiltonians, which is guaranteed by the condensable algebra being Lagrangian [90]). Since these examples and the construction of the stabilizer Hamiltonians will inform our subsequent analysis of mixed-state SPTs, we would encourage even SymTFT experts to peruse the remainder of this Section.
@@
Given the bulk SymTFT, one can, in principle, construct an effective lattice Hamiltonian for the boundary which is in correspondence with the set of condensed anyons on the dynamical boundary (see e.g., Refs. [91, 25, 29, 92, 93, 42].) We will not reconstruct these arguments here but, rather, will demonstrate how the anyon data of the SymTFT can be used to construct such a Hamiltonian. In the present case, since the bulk TO is given by $\mathcal{D}(\mathbb{Z}_{n})$, we choose the boundary degrees of freedom to be $n$-qudits, with generalized Pauli operators $\mathcal{X}_{j},\mathcal{Z}_{j}$ acting on each site. These operators satisfy the $\mathbb{Z}_{n}$ Pauli algebra, $\mathcal{X}_{j}^{n}=\mathcal{Z}_{j}^{n}=1$ and $\mathcal{Z}_{j}\mathcal{X}_{j}=e^{2\pi i/n}\mathcal{X}_{j}\mathcal{Z}_{j}$. As noted in Section II.1 and depicted in Fig. 2(b), the $\mathbb{Z}_{n}$ symmetry is generated by the magnetic (or $m$) Wilson line–on the edge, we can choose it to be generated by $\mathcal{X}\equiv\prod_{i}\mathcal{X}_{i}$. A boundary on which $\langle\mathcal{X}_{i}\rangle=+1$ would then be understood as one on which $m$ is condensed. Similarly, the $\mathbb{Z}_{n}$ order parameter is given by the $e$ line stretching from the reference boundary, where it is condensed, to the dynamical boundary. Since $e$ braids with a phase of $e^{2\pi i/n}$ with $m$, it is natural to associate the action of the $e$ line terminating on the dynamical boundary with the operator $\mathcal{Z}_{i}$, since the commutation relation $\mathcal{Z}_{i}\mathcal{X}=e^{2\pi i/n}\mathcal{X}\mathcal{Z}_{i}$ matches the bulk anyon braiding. A boundary corresponding to condensation of $e$ will thus exhibit long-range order in $\mathcal{Z}_{i}$; for instance, a state with $\langle\mathcal{Z}_{i}^{\dagger}\mathcal{Z}_{j}\rangle=1$ would correspond to such an $e$-condensed boundary. This latter operator may be understood as an $e$ line emanating from a point on the dynamical boundary and terminating at an adjacent point on the same boundary, indicating that an $e¯\overline{e}$ anyon pair can emerge from the bulk and condense on the boundary. This motivates the correspondence
@@
Exploiting this correspondence, we can write down a stabilizer Hamiltonian whose ground state subspace realizes a 1+11+1d gapped phase with the symmetry data encoded in the bulk SymTFT:
$H_{1d}=-\sum_{i}\left[(\mathcal{Z}_{i}^{\dagger}\mathcal{Z}_{i+1})^{k}+(\mathcal{X}_{i})^{n/k}+\text{ h.c.}\right]$ (10)
@@
Next, let us consider $G=\mathbb{Z}_{2}^{A}\times\mathbb{Z}_{2}^{B}$. The bulk $\mathcal{D}(G)$ TO admits a total of six gapped boundaries to vacuum, of which five correspond to picking subgroups of $G$ but only admit a trivial $2$-cocyle (these gapped boundaries thus all correspond to 1+11+1d phases where the $G$ symmetry is spontaneously broken). The remaining gapped boundary, which is obtained by picking the subgroup $H=\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ (thus there is no spontaneous symmetry breaking) along with a nontrivial cocycle, realizes the nontrivial $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SPT phase. The corresponding Lagrangian subgroup in this case is generated by
@@
as realizing the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ 1+11+1d SPT. This Hamiltonian is indeed the familiar cluster-state Hamiltonian [94, 95]. From Eq. (9), we see that the two terms in this Hamiltonian can be interpreted as corresponding to the condensation of $e_{A}m_{B}$ and $e_{B}m_{A}$, respectively, on the dynamical boundary of the bulk SymTFT. Note that, in the second term we chose $X_{i+1}^{A}$ rather than $X_{i}^{A}$ in order to ensure that the stabilizers commute; this choice is not dictated by the heuristic correspondence of Eq. (9). In the ground state, the stabilizers all have eigenvalue +1+1. The string order parameters characterizing this SPT phase are
@@
where the $\ldots$ denote products of $X_{A}$ ($X_{B}$) from sites $i$ to j1j-1 (sites $i+1$ to $j$) in the first (second) line.
@@
Finally, in anticipation of our analysis for mixed-state SPTs, we consider $G=\mathbb{Z}_{4}^{A}\times\mathbb{Z}_{4}^{B}$ and focus on a particular 1+11+1d phase exhibiting both symmetry breaking and SPT order under the residual symmetry. We choose $H=\mathbb{Z}_{4}^{AB}\times\mathbb{Z}_{2}^{A}$ to be the subgroup generated by the diagonal $\mathbb{Z}_{4}$ action and $\mathbb{Z}_{2}^{A}\subset\mathbb{Z}_{4}^{A}$ acting on the first factor [^8] This symmetry structure will reappear in our analysis of mixed-state examples, with the “ABAB” and “AA” symmetries corresponding, respectively, to weak and strong symmetries. See Sec. V.. We choose the Lagrangian subgroup
@@
$\begin{aligned} H_{1d}=-\sum_{i}\Big(&\mathcal{Z}_{A,i-1}^{2}\mathcal{X}_{A,i-1}^{2}\mathcal{Z}_{A,i-1}^{2}+\mathcal{Z}_{B,i-1}^{2}\mathcal{X}_{B,i-1}^{2}\mathcal{Z}_{B,i}^{2}\\ +&\mathcal{Z}^{2}_{A,i-1}\mathcal{Z}^{2}_{A,i}\mathcal{Z}^{2}_{B,i-1}\mathcal{Z}^{2}_{B,i}+\mathcal{X}^{2}_{A,i}\mathcal{X}^{2}_{B,i}\\ +&\mathcal{Z}^{\dagger}_{A,i-1}\mathcal{Z}_{A,i}\mathcal{X}_{A,i}\mathcal{Z}_{B,i-1}\mathcal{Z}^{\dagger}_{B,i}\mathcal{X}_{B,i}+\text{ h.c.}\Big)\end{aligned}.$ (16)
@@
$\begin{aligned} \langle\mathcal{Z}_{A,i}^{2}\mathcal{Z}_{B,i}^{2}\mathcal{Z}_{A,j}^{2}\mathcal{Z}_{B,j}^{2}\rangle&=1\\ \langle\mathcal{Z}_{A,i}^{2}\mathcal{X}_{A,i}^{2}\ldots\mathcal{X}_{A,j-1}^{2}\mathcal{Z}_{A,j}^{2}\rangle&=1\\ \langle\mathcal{Z}^{\dagger}_{A,i}\mathcal{Z}_{B,i}\mathcal{X}_{A,i+1}\mathcal{X}_{B,i+1}\ldots\mathcal{X}_{A,j}\mathcal{X}_{B,j}\mathcal{Z}_{A,j}\mathcal{Z}^{\dagger}_{B,j}\rangle&=1\end{aligned}.$ (17)